OpenStudy (anonymous):

45 is what percent of 15

OpenStudy (anonymous):

Let the percentage be: x (x/100)*15 = 45 15x/100 = 45 x = (45*100)/15 x = 3*100 Multiply and get the answer..

OpenStudy (anonymous):

300